# Constant Rate Scheduling: Constant-Rate Distributional Change for Efficient Training and Sampling in Diffusion Models

### v3

@@ 5.1 Results on Pixel-Space Diffusion Models
We trained the continuous-time diffusion models on LSUH Horse 256 $\times$ 256 and Bedroom 256 $\times$ 256 for unconditional image generation. Detailed experimental results are listed in Appendix F.1.
LSUN Horse 256 $\times$ 256: The evaluation results are summarized in Table 2. The rows shaded in gray represent the dependence on the training schedule. Although sample fidelity and mode coverage are generally considered to be in a trade-off relationship, CRS-$v_{x}$ achieved the highest performance for all metrics. This result indicates that CRS-$v_{x}$ is not merely balancing trade-offs between evaluation metrics but fundamentally improves the quality of the generated samples.
The unshaded rows represent the dependence on sampling schedules. CRS-$v_{x}+v_{\mathrm{FID}}$ achieved the highest performance for all metrics, indicating fundamental improvement in sample quality (see Appendix D for the hyperparameter setting of CRS-$v_{x}+v_{\mathrm{FID}}$). The FID score of 2.03 surpasses the current state-of-the-art FID score of 2.11 lsun_horse_sota .
LSUN Bedroom 256 $\times$ 256: We confirmed the advantage of CRS in the low-NFE regime by comparing it with the leading results reported in GITS noise_gits . We trained the continuous-time diffusion model and used CRS-$v_{x}$ for optimizing the training schedule. The results are listed in Table 2. The best FID scores at $\mathrm{NFE}=5$ and $\mathrm{NFE}=10$ were achieved when using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$ for optimizing the training and sampling schedules, respectively.
Table 1: Dependence on training and sampling schedules for LSUN Horse $256\times 256$ in pixel-space diffusion models at $\mathrm{NFE}=250$. SDE-DPM-Solver++(2M) was used as sampler. Bold values represent best training or sampling schedules for each metric.
@@
Table 2: FID scores evaluated on LSUN Bedroom 256 $\times$ 256 in pixel-space diffusion models at small NFEs. Bold values indicate best FID scores at each NFE. Results for rows other than "Ours" are taken from previously published study noise_gits .
@@
EDM UniPC sampler_unipc EDM 23.34 5.75
AMED-Solver sampler_amed 5.38
iPNDM sampler_ipndm GITS noise_gits 15.85 5.28
@@ 5.2 Results on Latent-Space Diffusion Models
We trained latent-space diffusion models using the pre-trained autoencoder VQ-4 used in latent diffusion models (LDMs) ldm . We used LSUN Church and Bedroom for unconditional image generation, and ImageNet for class-conditional image generation. The image resolution was set to $256\times 256$, which was reduced to $64\times 64$ in the latent space. We present the evaluation results on LSUN Church $256\times 256$. Detailed results are listed in Appendix F.2.
LSUN Church 256 $\times$ 256: The evaluation results are summarized in Table 4. The dependence on the training schedule is shown in the rows shaded in gray, and CRS-$v_{x}$ achieved the highest performance except for sFID. The unshaded rows represent the dependence on the sampling schedule, and CRS-$v_{x}$ further improves all evaluation metrics except for precision. These results indicate the superiority of CRS in latent-space diffusion models as well.
From the results in Table 4, CRS-$v_{x}$ appears to perform the best at both training and sampling; however, depending on the sampler and NFE, CRS-$v_{\varepsilon}$ is preferable at sampling.
5.3 Comparison with AYS Using EDM
Because $v_{x}(\alpha)$ is related to the KLUB introduced in AYS, we additionally compare CRS with AYS on FFHQ 64 $\times$ 64 using the pretrained model provided in EDMs edm . It is important to note that EDM adopts a variance-exploding process, while we have so far focused on the variance-preserving process. However, CRS can also be applied to the variance-exploding process by converting between $\alpha$ and the noise level in EDM $\sigma_{\mathrm{EDM}}$ via the log SNR, as shown below:
$\mathrm{logSNR}=\log\frac{\alpha^{2}}{1-\alpha^{2}}=\log\frac{1}{\sigma_{\mathrm{EDM}}^{2}}\Rightarrow\alpha=\frac{1}{\sqrt{1+\sigma_{\mathrm{EDM}}^{2}}}.$ (16)
The results are listed in Table 4. CRS-$v_{x}+v_{\mathrm{FID}}$ consistently outperformed AYS at all NFEs. Detailed results on pretrained EDM are listed in Appendix F.3.
Table 3: Dependence on training and sampling schedules for LSUN Church $256\times 256$ in latent-space diffusion models at $\mathrm{NFE}=30$. DPM-Solver++(2M) was used as sampler. Bold values represent best training or sampling schedules for each metric.
@@ 5.3 Comparison with AYS Using EDM
Table 4: FID scores evaluated on FFHQ 64 $\times$ 64 using the pretrained model provided in EDM. FID scores on AYS is taken from their original paper noise_ays .
Sampler Noise Schedule $\mathrm{NFE}=10$ $\mathrm{NFE}=20$ $\mathrm{NFE}=50$
DPM-Solver++(2M) AYS noise_ays 5.43 3.29 2.62
CRS-$v_{x}+v_{\mathrm{FID}}$ 4.70 2.87 2.52
@@ 5.4 Comparison of Noise Schedules
The CRS-based noise schedules in the pixel-space diffusion model are shown in Fig. 3. The hyperparameters of CRS-$v_{x}+v_{\mathrm{FID}}$ was set to $w_{x}=w_{\mathrm{FID}}=0.5$ and $\xi_{x}=\xi_{\mathrm{FID}}=1.0$. Since $v_{\mathrm{FID}}(\alpha)$ takes large values in regions where sample fidelity degrades in the diffusion process, CRS-$v_{\mathrm{FID}}$ allocates many timesteps to the region where $\alpha\simeq 1$. In contrast, CRS-$v_{x}$ allocates more timesteps to regions with small $\alpha$. This is due to the fact that the data prediction $\bm{x}_{\theta}(\bm{x},\alpha)$ can be expressed as $\bm{x}_{\theta}(\bm{x},\alpha)=\frac{\bm{x}-\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)}{\alpha}$ using the noise prediction, where $\alpha$ appears in the denominator. Thus, CRS-$v_{\mathrm{FID}}$ and CRS-$v_{x}$ exhibits complementary characteristics, and combining them yields an effective sampling schedule. Since the noise schedule based on CRS-$v_{\varepsilon}$ can be approximately reproduced by tuning the hyperparameters of CRS-$v_{x}+v_{\mathrm{FID}}$, we regard $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ as the fundamental compenents.
@@
Figure 3 shows the CRS-based noise schedules in the latent-space diffusion model. The appropriate noise schedules differ significantly between the pixel-space and lantet-space diffusion models. The most notable difference is that, in the latent-space diffusion models, fewer timesteps are allocated to the region where $\alpha\simeq 1$. This can be attributed to the fact that, in latent-space diffusion models, the perceptual expression is largely generated by the autoencoder, and the autoencoder can absorb minor deviations in the generated latent variables. In contrast, the pixel-space diffusion models directly generate images, so careful control of the noise schedule is crucial at $\alpha\simeq 1$. We did not evaluate any combination of multiple distance metrics in latent-space diffusion models, because all the CRS-based noise schedules have a similar functional form.
@@
Figure 2: Noise schedules on LSUN Horse 256 $\times$ 256 in pixel-space diffusion model.
@@
Figure 3: Noise schedules on LSUN Church 256 $\times$ 256 in latent-space diffusion model.
@@ 6 Limitations
This paper focuses on developing CRS and validating its effectiveness under fundamental experimental settings. However, to further strengthen the generality of CRS, additional investigations are needed from the following perspectives: 1) generality across various generation conditions, 2) scalability of the distance metrics with respect to dataset size and data dimensionality, 3) effectiveness in domains beyond image generation. While $v_{\mathrm{FID}}(\alpha)$ proved effective as the sampling schedule in pixel-space diffusion models, it is a distance metric specialized for image generation. Therefore, when applying CRS to other domains, it may be necessary to introduce new distance metrics that are better suited to the characteristics of the target domain.

### v4

@@ 5.2 Results on Pixel-Space Diffusion Models
We trained continuous-time diffusion models on LSUN Horse $\times256$ and LSUN Bedroom $\times256$ for unconditional image generation. Detailed experimental results are provided in Appendix F.1.
LSUN Horse $\times256$: The evaluation results are summarized in Table 2. Rows shaded in gray represent results obtained under different training schedules. Although sample fidelity and mode coverage are generally considered to be in a trade-off relationship, CRS-$v_{x}$ achieved the highest scores across all evaluation metrics. This indicates that CRS-$v_{x}$ does not merely balance fidelity and diversity but leads to an overall improvement in sample quality.
The unshaded rows represent the dependence on sampling schedules. CRS-$v_{x}+v_{\mathrm{FID}}$ outperformed all baselines across all metrics, indicating a fundamental improvement in generation quality. Notably, our model achieved an FID score of 2.03, surpassing the previous state-of-the-art score of 2.11 reported in Kumari et al. (2022).
LSUN Bedroom $\times256$: We further validated the effectiveness of CRS in low-NFE regimes by comparing it with strong baselines reported in GITS (Chen et al., 2024). We trained continuous-time diffusion models with CRS-$v_{x}$ used for optimizing the training schedule. Table 2 lists the results. At $\mathrm{NFE}=5$ and $\mathrm{NFE}=10$, the best FID scores were achieved by using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$ for optimizing training and sampling schedules, respectively.
As shown in Table 20 in Appendix F.1, we also tuned the $\rho$ parameter in the EDM sampling schedule. Even with the optimal $\rho$ setting, CRS-$v_{x}+v_{\mathrm{FID}}$ consistently outperformed EDM.
Table 1: Performance of pixel-space diffusion models on LSUN Horse $\times256$ with $\mathrm{NFE}=250$, using SDE-DPM-Solver++(2M) as the sampler. Gray-shaded rows indicate variations in the training schedule, while unshaded rows correspond to different sampling schedules. Bold values highlight the best performance for each metric under varying training or sampling schedules.
@@
Table 2: FID scores for LSUN Bedroom $\times256$ in pixel-space diffusion models under low NFE conditions. Our results are compared with prior work from Chen et al. (2024). Bold entries indicate the best FID achieved at each NFE setting.
@@
EDM UniPC (Zhao et al., 2023) EDM 23.34 5.75
AMED-Solver (Zhou et al., 2024) 5.38
iPNDM (Zhang and Chen, 2023) GITS (Chen et al., 2024) 15.85 5.28
@@ 5.3 Results on Latent-Space Diffusion Models
We trained latent-space diffusion models using the pre-trained autoencoder from LDMs (Rombach et al., 2022). In our experiments, we adopted the VQ-regularized variant for all datasets, which is the configuration primarily used in the publicly available pretrained models on GitHub (https://github.com/CompVis/latent-diffusion). Although a downsampling factor of $f=8$ is also commonly used, we set $f=4$ (VQ-f4), as it is expected to yield better FID scores according to Table 10 in the LDMs paper. LSUN Church and LSUN Bedroom were used for unconditional image generation, and ImageNet was used for class-conditional image generation. All images were resized to $\times256$ and encoded into $\times64$ latent representations. Here, we present the evaluation results on LSUN Church $\times256$. Detailed results are provided in Appendix F.2.
LSUN Church $\times256$: Table 4 summarizes the results. Rows shaded in gray show the effect of different training schedules, where CRS-$v_{x}$ achieved the best performance on all metrics except sFID. Unshaded rows show the dependence on the sampling schedule, and again, CRS-$v_{x}$ achieved the best performance on all metrics except precision.
These results indicate that CRS is also effective in latent-space diffusion models. Although CRS-$v_{x}$ performs best in most settings, we observe that depending on the sampler and NFEs, CRS-$v_{\varepsilon}$ can be preferable for sampling.
Table 3: Performance of latent-space diffusion models on LSUN Church $\times256$ with $\mathrm{NFE}=30$, using DPM-Solver++(2M) as the sampler. Gray-shaded rows indicate variations in the training schedule, while unshaded rows correspond to different sampling schedules. Bold values highlight the best performance for each metric under varying training or sampling schedules.
@@
Table 4: FID scores on CIFAR10 $\times32$ and FFHQ $\times64$ evaluated using the pretrained model provided by EDM (Karras et al., 2022). Bold numbers indicate the best sampling schedule for each combination of dataset, sampler, and NFE. For reference, the FID scores of AYS (Sabour et al., 2024) and LD3 (Tong et al., 2025) are taken directly from their original papers.
Sampling CIFAR10 $\times32$ FFHQ $\times64$
Sampler schedule $\mathrm{NFE}=10$ $\mathrm{NFE}=20$ $\mathrm{NFE}=10$ $\mathrm{NFE}=20$
DPM-Solver++(2M) AYS 2.98 2.10 5.43 3.29
GITS 4.03 2.32 5.51 3.81
LD3 3.38 2.36 3.98 2.89
CRS-$v_{x}+v_{\mathrm{FID}}$ 3.92 2.27 5.35 3.10
UniPC(3M) GITS 3.50 1.99 4.52 2.71
LD3 2.84 3.27
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.52 1.99 5.23 2.64
iPNDM(3M) GITS 2.73 2.08 3.96 3.09
LD3 2.38 3.25
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.87 2.11 4.07 2.75
iPNDM(4M) GITS 2.49 2.02 3.62 3.00
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.46 2.02 3.45 2.53
5.4 Comparison with Recent Sampling Schedules
Using the pretrained models released with EDM (Karras et al., 2022), we compare the performance of CRS as a sampling schedule against AYS (Sabour et al., 2024), GITS (Chen et al., 2024), and learning to discretize denoising diffusion ODEs (LD3) (Tong et al., 2025). All evaluations in this subsection are conducted using the official repository provided in the GITS paper, ensuring full consistency with their experimental protocol.
It is important to note that EDM adopts a variance-exploding formulation, whereas all of our previous experiments have been conducted under the variance-preserving setting. However, CRS can still be applied to variance-exploding processes by converting between $\alpha$ and the EDM noise level $\sigma_{\mathrm{EDM}}$ via the log-SNR, as shown below:
$\log\mathrm{SNR}=\log\left(\frac{\alpha^{2}}{1-\alpha^{2}}\right)=\log\left(\frac{1}{\sigma_{\mathrm{EDM}}^{2}}\right)\quad\Rightarrow\quad\alpha=\frac{1}{\sqrt{1+\sigma_{\mathrm{EDM}}^{2}}}.$ (17)
We also note that, whereas all earlier experiments in this paper used the noise-prediction parameterization, EDM adopts the data-prediction parameterization with preconditioning.
The evaluation results are presented in Table 4. The optimal sampling schedule varies across datasets, samplers, and NFE configurations. On average, LD3 achieves the highest performance, while CRS consistently outperforms GITS and can therefore be regarded as the second-best sampling schedule overall. Although CRS underperforms LD3 as a sampling schedule, a key advantage of CRS is that it also provides a unified framework for optimizing the training schedule, which LD3 does not address. Further improving the performance of CRS will likely require identifying more effective discrepancy measures, and exploring such alternatives is an important direction for future work.
For completeness and fairness, we note that CRS-$v_{x}+v_{\mathrm{FID}}$ requires both the computation of $v_{\mathrm{FID}}(\alpha)$ and hyperparameter tuning, and the time needed to optimize the sampling schedule is substantially larger than that of GITS or LD3. However, sampling-schedule optimization is performed only once after training the diffusion model. Therefore, as long as the tuning cost remains small relative to the total training time, we believe that performance under the best hyperparameter configuration is the most meaningful criterion for comparison.
@@ 5.5 Comparison of Noise Schedules
Figure 3 shows examples of CRS-based noise schedules in pixel-space diffusion models, computed using different discrepancy measures. For CRS-$v_{x}+v_{\mathrm{FID}}$, the hyperparameters were set to $w_{x}=w_{\mathrm{FID}}=0.5$ and $\xi_{x}=\xi_{\mathrm{FID}}=1.0$.
@@
Since $v_{\mathrm{FID}}(\alpha)$ tends to be large in regions where sample fidelity degrades (i.e., near $\alpha\simeq 1$), CRS-$v_{\mathrm{FID}}$ allocates more timesteps to these regions. In contrast, CRS-$v_{x}$ allocates more timesteps to regions with smaller $\alpha$. This is due to the formulation of the data prediction $\bm{x}_{\theta}(\bm{x},\alpha)=\frac{\bm{x}-\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)}{\alpha}$, where $\alpha$ appears in the denominator, making the data prediction more sensitive to small changes in $\alpha$ when $\alpha$ is small. Thus, CRS-$v_{\mathrm{FID}}$ and CRS-$v_{x}$ exhibit complementary behaviors, and combining both yields an effective and balanced sampling schedule.
Furthermore, we observed that the noise schedule derived from CRS-$v_{\varepsilon}$ can be approximately reproduced by appropriately tuning the hyperparameters of CRS-$v_{x}+v_{\mathrm{FID}}$. Therefore, in practice, $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ serve as the two principal components for constructing robust sampling schedules.
@@
Figure 3 illustrates CRS-based noise schedules for latent-space diffusion models. The optimized schedules differ notably between pixel-space and latent-space models. In latent-space diffusion, fewer timesteps are allocated to the region where $\alpha\simeq 1$. This difference can be attributed to the nature of latent-space diffusion models: perceptual expression is largely reconstructed by the autoencoder, and the autoencoder is robust to minor deviations in the generated latent variables. Consequently, precise control near $\alpha\simeq 1$ is less critical. In contrast, pixel-space models directly generate the final image, so the schedule must allocate sufficient timesteps near $\alpha\simeq 1$ to preserve high-fidelity details.
Because all discrepancy measures result in similar noise schedules for latent-space models, we did not evaluate combinations of multiple discrepancy measures.
@@
Figure 2: Noise schedules computed using different discrepancy measures on LSUN Horse $\times256$ in the pixel-space diffusion model.
@@
Figure 3: Noise schedules computed using different discrepancy measures on LSUN Church $\times256$ in the latent-space diffusion model.
5.6 Computational Overhead of CRS
CRS requires computing $v(\alpha)$. Below, we report the computational overhead in scenarios where the cost is expected to be most significant.
Cost of computing $v_{\mathrm{FID}}(\alpha)$: Although $v_{\mathrm{FID}}(\alpha)$ needs to be computed only once before training the diffusion model, its computation requires processing the entire training set to compute feature statistics, and therefore the cost scales linearly with the number of images.
Table 6 reports both the cost relative to a single training epoch and the cost relative to the total training time. Because the computation of $v_{\mathrm{FID}}(\alpha)$ and the duration of one training epoch both scale proportionally with dataset size, the cost relative to one training epoch remains nearly constant across datasets for both pixel-space and latent-space diffusion models. The relative cost appears larger for latent-space diffusion models simply because one latent-space epoch is much faster than one pixel-space epoch. When compared to the total training time, the overhead of computing $v_{\mathrm{FID}}(\alpha)$ remains below 15% even in the worst case.
Overhead of computing $v_{x}(\alpha)$ during training: Table 6 compares the training time when using the linear versus CRS-$v_{x}$ as the training schedule. Because CRS-$v_{x}$ requires evaluating $\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})$ during training, the overall training time increases by approximately 20–30%. Nevertheless, CRS-$v_{x}$ typically accelerates the convergence of FID scores during training (see Figs. 13 and 13 in Appendix F for details), which can largely compensate for the per-epoch overhead.
Table 5: Wall-clock computational cost of computing $v_{\mathrm{FID}}(\alpha)$. We report both the cost relative to one training epoch and the cost relative to the total training time for each diffusion setting.
Computational cost
Diffusion type Dataset Relative to one training epoch Relative to total training time
Pixel space LSUN Horse $\sim$5 epochs $\sim$10%
LSUN Bedroom $\sim$5 epochs $\sim$10%
Latent space LSUN Church $\sim$14 epochs $\sim$1.4%
LSUN Bedroom $\sim$14 epochs $\sim$14%
ImageNet $\sim$14 epochs $\sim$2.8%
Table 6: Comparison of wall-clock training time when using linear and CRS-$v_{x}$ as the training schedule. CRS-$v_{x}$ requires evaluating $\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})$ during training, resulting in a 20–-30% increase in overall training time.
Diffusion type Dataset Linear CRS-$v_{x}$ Increase
Pixel space LSUN Horse 170 hours 215 hours +26.5%
Latent space LSUN Church 85 hours 105 hours +23.5%
LSUN Bedroom 190 hours 245 hours +28.9%
ImageNet 430 hours 525 hours +22.1%
5.7 A Practical Alternative to FID-Based Scheduling
In pixel-space diffusion models, CRS-$v_{x}+v_{\mathrm{FID}}$ consistently achieves the best performance as a sampling schedule. However, for large-scale training datasets, computing $v_{\mathrm{FID}}(\alpha)$ becomes prohibitively expensive.
As shown in the subsection 5.5, CRS-$v_{\mathrm{FID}}$ allocates a large number of timesteps to regions near $\alpha\simeq 1$, resulting in a schedule shape similar to that of the cosine schedule. Motivated by this observation, we evaluate a lightweight alternative: CRS-$v_{x}+v_{\mathrm{cos}}$, which combines CRS-$v_{x}$ with the cosine schedule (see Eq. 13 for $v(\alpha)$ of the cosine schedule).
Table 7 compares the FID scores of CRS-$v_{x}+v_{\mathrm{FID}}$ and CRS-$v_{x}+v_{\mathrm{cos}}$ on LSUN Bedroom $\times256$. In all settings, CRS-$v_{x}+v_{\mathrm{cos}}$ achieves FID scores comparable to CRS-$v_{x}+v_{\mathrm{FID}}$, demonstrating that CRS-$v_{x}+v_{\mathrm{cos}}$ is a practical and effective alternative when $v_{\mathrm{FID}}(\alpha)$ is costly to compute.
Table 7: Evaluation of CRS-$v_{x}+v_{\mathrm{cos}}$ as a practical alternative to CRS-$v_{x}+v_{\mathrm{FID}}$ for sampling in pixel-space diffusion models on LSUN Bedroom $\times256$. Despite not requiring the computation of $v_{\mathrm{FID}}(\alpha)$, CRS-$v_{x}+v_{\mathrm{cos}}$ achieves similar FID scores.
Sampler Sampling Schedule $\mathrm{NFE}=5$ $\mathrm{NFE}=10$
DPM-Solver++(2M) CRS-$v_{x}+v_{\mathrm{FID}}$ 14.02 4.88
CRS-$v_{x}+v_{\mathrm{cos}}$ 14.10 4.46
UniPC CRS-$v_{x}+v_{\mathrm{FID}}$ 21.42 3.30
CRS-$v_{x}+v_{\mathrm{cos}}$ 20.77 3.30
@@ 6 Limitations
This paper focuses on developing CRS and empirically validating its effectiveness under fundamental experimental settings. While the results are promising, there are several directions for future work to further enhance the generality, applicability, and performance of CRS. Below, we summarize six key limitations and avenues for future work.
(1) Theoretical analysis.
The goal of this work is to provide a theoretically motivated and empirically effective framework capable of incorporating a wide range of discrepancy measures. However, we do not provide a theoretical guarantee that minimizing the maximum discrepancy necessarily leads to improved sample quality. Establishing a formal connection between distributional change and sample quality remains an important direction for future research.
(2) Evaluation on more complex tasks.
Our experiments focus on fundamental settings—unconditional and class-conditional image generation. We did not evaluate CRS in more complex practical scenarios such as text-conditional generation or classifier-free guidance (CFG). CFG is known to improve FID at the cost of reduced recall, and an adaptive noise schedule may help mitigate this trade-off. Evaluating CRS together with CFG or other conditioning mechanisms could further reveal the benefits of schedule optimization.
(3) Applicability beyond image generation.
Although our experiments are limited to image generation, CRS is derived from general diffusion-model principles and is therefore expected to generalize to other domains. In this work, $v_{\mathrm{FID}}(\alpha)$ proved effective as a sampling discrepancy measure for pixel-space diffusion models; however, it is specific to image generation. Applying CRS to non-image modalities will require identifying discrepancy measures that reflect the characteristics of each target domain.
(4) Improving discrepancy measures.
As a sampling schedule optimizer, CRS currently performs slightly worse than LD3, suggesting that further improvements may be achieved by designing better discrepancy measures. A promising direction is to build on advances in schedule-optimization theory such as AYS. As shown in Appendix B, CRS-$v_{x}$ can be interpreted as employing a simplified variant of the KLUB introduced in AYS. Thus, future theoretical progress in schedule optimization may naturally lead to new discrepancy measures compatible with CRS.
With respect to image-specific metrics, although our experiments primarily relied on FID, many feature-based perceptual metrics have been proposed for image generation. For example, LPIPS (Zhang et al., 2018) is a strong candidate for capturing visually meaningful differences. Since FID and LPIPS depend on feature extractors, a key challenge is designing robust feature representations for noised data, which are essential for accurately assessing distributional change. Self-supervised learning on noised samples is a promising direction, as it requires no labels and can be made noise-level aware (e.g., by predicting or conditioning on $\alpha$). Such learned embeddings may lead to more expressive and domain-adaptive discrepancy measures.
(5) Generality across prediction parameterizations and architectures.
Our experiments primarily evaluate CRS under the noise-prediction parameterization. However, the data- and velocity-prediction formulations are also widely used in recent diffusion models, and the generality of CRS under these parameterizations has not been thoroughly assessed.
Furthermore, modern diffusion architectures such as Diffusion Transformers (Peebles and Xie, 2023) were not evaluated. Because CRS is grounded in the dynamics of distributional evolution rather than in model architecture, we expect architectural dependence to be limited, but empirical validation remains an important direction for future work.
(6) Scalability to dataset size and image resolution.
Although CRS-$v_{x}+v_{\mathrm{FID}}$ is effective for optimizing sampling schedules in pixel-space diffusion models, it requires both the computation of $v_{\mathrm{FID}}(\alpha)$ and hyperparameter tuning. In particular, the computation of $v_{\mathrm{FID}}(\alpha)$ scales linearly with the number of training images, making this variant impractical for large-scale datasets. Although Section 5.7 demonstrates that CRS-$v_{x}+v_{\mathrm{cos}}$ can serve as a practical alternative, fundamentally improving scalability will require developing new discrepancy measures that are both computationally efficient and robust to noise-level variation.
Furthermore, all of our experiments were conducted at a resolution of $\times256$. Validating whether CRS remains effective for higher-resolution image generation, and analyzing how the optimal noise schedule depends on image size, are compelling directions for future research.